How to solve the equation?
Log base 2 (x+2) – log base 2 (x-5) = 3

Answer by stanbon(75887)

solve the equation?
Log base 2 (x+2) – log base 2 (x-5) = 3
-----
log[(x+2)/(x-5)] = 3
------
(x+2)/(x-5) = 10^3
------
x+2 = 1000x - 5000
-----
999x = -4998
x = -5.003
--------
Cheers,
Stan H.
--------------